# Towards Proof Repair in Cubical Agda

### 02 Feb 2024 v2

@@ 3.2. Proof Repair in Cubical Agda
We make one simplifying assumption to accommodate these set quotients. For a type T defined using set quotients, we require that the motive provided to the eliminator, (P: T normal-$\to$ Set) satisfies the condition (x: T) normal-$\to$ isSet (P x). This condition states that, for any x : T, the type P x satisfies the uniqueness of identity proofs, or, in the language of Cubical Agda, is an h-set. Because we forbid higher inductive types aside from set quotients, this condition will be satisfied for many applications in software engineering and verification.
@@
3.3. Correctness
We would like to be able to prove that our function and proofs were repaired correctly. Stating what it means for repair to have occurred correctly is nontrivial. Intuitively, we would like for terms on the new type to behave the same as terms on the old type. However, because our old and new types are different, we cannot simply state this as a homogeneous equality. Instead, we want to know that the terms behave the same up to the equivalence we repair across.
This was done metatheoretically in a univalent type theory for the Pumpkin Pi transformation that we build off of, but it could not be done internally in general, and it was not proven for any particular example at all. In contrast, in Cubical Agda, we have access to heterogeneous equalities in the form of PathP types. This gives us the power needed to formalize the correctness theorems about repair from the Pumpkin Pi paper in Cubical Agda, and even to prove them correct on an example type for the first time.
@@ 3.3. Correctness
  lamOK: {T} {F}
    (f: (t: T i0) normal-$\to$ F i0 t) (f’: (t: T i1)normal-$\to$ F i1 t)
@@
    PathP ($\lambda$ i normal-$\to$ $\forall$ (t : T i) normal-$\to$ F i t) f f
@@
Other rules are stated specifying the types being repaired. For example, we proved that the repaired eliminator we defined for a simple quotient equivalence was correct. Our source type was $\mathbb{N}$, and our target was Int / rInt, where Int = $\mathbb{N}\uplus\mathbb{N}$ and rInt is the relation placing inl n and inr n in the same equivalence class. The dependent constructors and eliminators correspond to the usual ones for $\mathbb{N}$. The correctness condition for a repaired eliminator for a given configuration was stated externally in the Pumpkin Pi paper, but it was not proven for any type. We adapted this theorem to Cubical Agda for our example (Figure 4) and, for the first time, proved that it held (see supplementary material).
Even when one can internally prove individual rules of the transformation correct for a particular example change in types, it can be prohibitively challenging to actually compose those proofs to get proofs of the correctness of particular repaired proofs. We discuss the reasons for this in Section 5.
@@ 4. Case Studies
This paper centers two case studies for manual proof repair in Cubical Agda. First, we conduct repair between two representations of the integers (Section 4.1). Second, we study two common implementations of the queue data structure, and how we can repair from one to the other (Section 4.2). Both of these case studies are empowered by the presence of quotient types in Cubical Agda.
@@ 4.1. Adding, Fast and Slow
Our first case study is mathematically motivated. We consider a change in the type representing the integers. We repair addition and proofs about addition from one representation to the other. Finally, we recover the repaired proofs for a more efficient version of addition over the repaired type.
@@
We start with the default implementation of the integers in Cubical Agda—two copies of $\mathbb{N}$ glued together with one of them reversed, which we call $\mathbb{Z}$.
@@
Figure 5. The types of our integer representations.
@@
We will repair functions and proofs about $\mathbb{Z}$ to use a representation of integers that may be more familiar to set theorists: Instead of two copies of $\mathbb{N}$ glued together at the ends, we consider the integers as elements of $\mathbb{N}\times\mathbb{N}/\sim$, where $(x_{1},x_{2})\sim(y_{1},y_{2})\iff x_{1}+y_{2}=x_{2}+y_{1}$. [^6] We can view this an an instance of constructing the Grothendieck group from the commutative monoid $\mathbb{N}$, and hence the integers arise as the unique group satisfying the universal property that any monoid homomorphism out of $\mathbb{N}$ can be uniquely extended to a group homomorphism out of $\mathbb{Z}$. We would not be able to express this naturally in Coq, but we can represent this in Cubical Agda by defining this relation explicitly and then quotienting $\mathbb{N}\times\mathbb{N}$ by that relation. We refer to the resulting quotient type as GZ. The definitions of $\mathbb{Z}$ and GZ can be found in Figure 5. In order to do repair between these two types, we need them to be isomorphic. In this case, the isomorphism is the map we expect: pos n maps to [(n , 0)], while negsuc n maps to [(0 , n + 1)]. Verifying that this map is bijective is straightforward.
$\mathbb{Z}$Pos : $\mathbb{N}$ $\to \mathbb{Z}$
$\mathbb{Z}$NegSuc : $\mathbb{N}$ $\to \mathbb{Z}$
$\mathbb{Z}$ : (P : $\mathbb{Z}$ normal-$\to$ Type) normal-$\to$
  ((n : $\mathbb{N}$) normal-$\to$ P ($\mathbb{Z}$Pos n)) normal-$\to$
  ((n : $\mathbb{N}$) normal-$\to$ P ($\mathbb{Z}$NegSuc n)) normal-$\to$
  (x : $\mathbb{Z}$) normal-$\to$ P x
$\iota\mathbb{Z}$Pos : (P : $\mathbb{Z}$ normal-$\to$ Set) normal-$\to$
  (posP : (n : $\mathbb{N})$ normal-$\to$ P ($\mathbb{Z}$Pos n)) normal-$\to$
  (negSucP : (n : $\mathbb{N})$ normal-$\to$ P ($\mathbb{Z}$NegSuc n)) normal-$\to$
  (n : $\mathbb{N}$) normal-$\to$
  (Q : P ($\mathbb{Z}$Pos normal-$\to$ Set) normal-$\to$
  Q ($\mathbb{Z}$ P posP negSucP ($\mathbb{Z}$Pos n)) normal-$\to$
  Q (posP n)
$\iota\mathbb{Z}$NegSuc : (P : $\mathbb{Z}$ normal-$\to$ Set) normal-$\to$
  (posP : (n : $\mathbb{N}$) normal-$\to$ P ($\mathbb{Z}$Pos n)) normal-$\to$
  (negSucP : (n : $\mathbb{N})$ normal-$\to$ P ($\mathbb{Z}$NegSuc n)) normal-$\to$
  (n : $\mathbb{N}$) normal-$\to$
  (Q : P ($\mathbb{Z}$NegSuc normal-$\to$ Set) normal-$\to$
  Q ($\mathbb{Z}$ P posP negSucP ($\mathbb{Z}$NegSuc n)) normal-$\to$
  Q (negSucP n)
depConstrGZPos : $\mathbb{N}$ normal-$\to$ GZ
depConstrGZNegSuc : $\mathbb{N}$ normal-$\to$ GZ
depElimGZ : (P : GZ normal-$\to$ Type) normal-$\to$
  ((x : GZ) normal-$\to$ isSet (P x)) normal-$\to$
  ((n : $\mathbb{N})$ normal-$\to$ P (depConstrGZPos n)) normal-$\to$
  ((n : $\mathbb{N})$ normal-$\to$ P (depConstrGZNegSuc n)) normal-$\to$
  (x : GZ) normal-$\to$ P x
$\iota$GZPos : (P : GZ normal-$\to$ Set) normal-$\to$
  ((x : GZ) normal-$\to$ isSet (P x)) normal-$\to$
  (posP : (n : $\mathbb{N})$ normal-$\to$ P (depConstrGZPos n)) normal-$\to$
  (negSucP : (n : $\mathbb{N})$ normal-$\to$ P (depConstrGZNegSuc n)) normal-$\to$
  (n : $\mathbb{N}$) normal-$\to$
  (Q : P (depConstrGZPos normal-$\to$ Set) normal-$\to$
  Q (depElimGZ P posP negSucP (depConstrGZPos n)) normal-$\to$
  Q (posP n)
$\iota$GZNegSuc : (P : GZ normal-$\to$ Set) normal-$\to$
  ((x : GZ) normal-$\to$ isSet (P x)) normal-$\to$
  (posP : (n : $\mathbb{N}$) normal-$\to$ P (depConstrGZPos n)) normal-$\to$
  (negSucP : (n : $\mathbb{N})$ normal-$\to$ P (depConstrGZNegSuc n)) normal-$\to$
  (n : $\mathbb{N}$) normal-$\to$
  (Q : P (depConstrGZNegSuc normal-$\to$ Set) normal-$\to$
  Q (depElimGZ P posP negSucP (depConstrGZNegSuc n)) normal-$\to$
  Q (negSucP n)
Figure 6. Configuration for our integer representations. Reversed versions of the $\iota$ rules, denoted with a “${}^{-}$”, are omitted.
@@
To repair functions and proofs across this change, we first must decompose our isomorphism into a configuration, which we do in Figure 6. Here, we can see the dependent constructors, eliminators, and $\iota$-reduction rules for both types. In addition to the listed $\iota$-reduction rules, there are also reversed versions, which we omit for the sake of space.
Notice that the types for the components indeed correspond with each other, and both sides of the configuration share the same inductive structure. The only change in types here new to Cubical Agda compared to Pumpkin Pi is that every motive P : GZ normal-$\to$ Set comes with the requirement that ((x : GZ) normal-$\to$ isSet (P x)).
Thus, with the configuration defined, we can perform repair according to the procedure discussed in Section 3. To give one concrete example of a term being repaired, we can consider the case of repairing addition from $\mathbb{Z}$ to GZ. First, we give the standard definition of addition in $\mathbb{Z}$:
@@
Then, we can follow the repair algorithm to get the term:

### 15 Mar 2024 v3

@@ 4.1. Quotient Types: Internal
For example, the type $\mathbb{N}/2$ from Section 2.1 can be captured with the following type:
  ~ : $\mathbb{N}$ $\to \mathbb{N}$ normal-$\to$ Type
  ~ n m = (n mod 2) $\equiv$ (m mod 2)
  $\mathbb{N}/2$ = $\mathbb{N}$ / $\sim$
The proof of equality between [ 0 ] and [ 2 ] is written eq/ [ 0 ] [ 2 ] refl.
Since quotient types are internal, a quotient type equivalence in Cubical Agda is simply a type equivalence that happens to be between quotient types. For the type above, we can construct an equivalence with the booleans using isoToEquiv: Iso A B normal-$\to$ A $\simeq$ B which allows us to get from any isomorphism to an equivalence. Here, we can apply isoToEquiv to the isomorphism between $\mathbb{N}/2$ and the booleans, yielding a type equivalence.
4.2. Transformation: Straightforward
[…]
We make one simplifying assumption to accommodate these set quotients. For a type T defined using set quotients, we require that the motive provided to the eliminator, (P: T normal-$\to$ Set), satisfies the condition (x: T) normal-$\to$ isSet (P x). This condition states that, for any x : T, the type P x satisfies the uniqueness of identity proofs, or, in the language of Cubical Agda, is an h-set. Because we forbid higher inductive types aside from set quotients, this condition will be satisfied for many applications in software engineering and verification.
@@ 4.2. Transformation: Straightforward
4.3. Correctness: Some Internal
We would like to be able to prove that our functions and proofs were repaired correctly. Intuitively, we would like for terms on the new type to behave the same as terms on the old type. However, because our old and new types are different, we cannot simply state this as a homogeneous equality. Instead, we want to know that the terms behave the same up to the equivalence we repair across.
This was done metatheoretically in a univalent type theory for the Pumpkin Pi transformation that we build off of, but it could not be done internally in general, and it was not proven for any particular example at all. In contrast, in Cubical Agda, we have access to heterogeneous path equalities in the form of PathP types. This gives us the power needed to formalize the correctness theorems about repair from the Pumpkin Pi paper in Cubical Agda, and even to prove them correct on an example type equivalence for the first time.
[…]
    PathP ($\lambda$ i normal-$\to$ $\forall$ (t : T i) normal-$\to$ F i t) f f
[…]
@@ 4.3. Correctness: Some Internal
Other rules are stated specifying the types being repaired. For example, we proved that the repaired eliminator we defined for a simple quotient equivalence was correct. Our source type was $\mathbb{N}$, and our target was Int / rInt, where Int = $\mathbb{N}\uplus\mathbb{N}$ and rInt is the relation placing inl n and inr n in the same equivalence class. The dependent constructors and eliminators correspond to the usual ones for $\mathbb{N}$. The correctness condition for a repaired eliminator for a given configuration was stated externally in the Pumpkin Pi paper, but it was not proven for any type. We adapted this theorem to Cubical Agda for our example (Figure 4) and, for the first time, proved that it held (see equivalence_int_abs.agda in the supplementary material).
One thing to note is that, even when we can internally prove individual rules of the transformation correct, it can be prohibitively challenging to actually compose those proofs to get proofs of the correctness of particular repaired proofs. Our correctness theorem is stated in terms of PathPs, but showing that two PathPs are equal requires reasoning about the equality of equality proofs between types. This means leaving the h-set fragment of cubical type theory (imposed by squash/) and entering the fragment that is proof relevant: identity proofs need not be unique, and specific proofs may be needed for goals. If our type equality proof can be written as $\lambda$ i normal-$\to$ Q (p i) where p is a path in an h-set and Q is a dependent product on elements of that h-set, we can use uniqueness of identity proofs to rewrite p to any other path with the same endpoints. But it is not always simple, or perhaps even possible, to frame the equality proof in this way.
Concretely, the way this often manifested is that we were able to compose the correctness proofs to show the correctness of repaired functions, like addition:
  addCorrect : $\forall$ (a b : $\mathbb{N}$) (a b : Int / rInt) normal-$\to$
@@
But we could not do the same for many more interesting proofs. For example, we tried to prove that a proof of addition being commutative was repaired correctly. Trying to compose the correctness proofs of the transformation rules in the same order as the transformation rules themselves at the type level resulted in a goal of a PathP along the equality:
  $\lambda$ i normal-$\to$ addCommCorrectType
    zero depConstrInt/rInt0 depConstr0Correct
@@
    i
But composing the correctness proofs in the natural order at the term level to attempt to fill that hole produced a PathP along the equality:
  $\lambda$ i normal-$\to$
    depConstrSCorrect i
@@
    depConstrSCorrect i
[…]
We failed to reconcile these two equality types. This is why we say that we were able to show just some correctness proofs internally, even though in theory, it should be possible to show all of them. Fully reckoning with proof relevance so that our correctness proofs compose correctly requires additional work. For this reason, we have yet to implement internal correctness proofs in the case studies we discuss in Section 6.
4.4. Automation: None
All of our proof repair work in Cubical Agda is completely manual, even if algorithmic. This highlights a major cost of working in Cubical Agda, despite the many positive type-theoretic properties that allow us to represent quotient types natively and correctness proofs internally. Standing in the way of automation for proof repair across quotient type equivalences in Cubical Agda are both engineering and theroetical limitations. The engineering challenges include:
  1. (1)
  2. (2)
    No metalanguage, no side effects. In Pumpkin Pi, side effects in the metalanguage of Coq were essential for the implementation of persistent caching and other features that were necessary to build practical, performant automation. Cubical Agda does not expose a metalanguage.
  3. (3)
    Lack of documentation and examples. There are a handful of domain-specific reflective tactics currently available in the standard library, such as NatSolver and CommRingSolver. Beyond these narrow domains, there is little documentation on how to safely build more general automation in a manner that does not increase the trusted computing base, especially when it comes to interacting with the type checker monad or any of Cubical Agda’s internals.
There are also significant theoretical challenges to resolve:
  1. (1)
    Higher equalities. We do not know how to adapt automated proof repair techniques to track and discharge proof obligations unique to cubical type theory, like the interval type and boundary conditions. In particular, we do not know of a technique or decision procedure that can reason about general path equalities and higher inductive types.
  2. (2)
    Proof relevance. Cubical type theory is proof relevant. A proof relevant type theory requires one to care about the manner in which a goal was proved, a detail that poses significant challenges to building automation in Cubical Agda. Personal correspondence with Cubical Agda developers indicates hesitance to build automation infrastructure due to theoretical concerns about automation and proof relevance.
5. External View: Proof Repair with Setoids in Coq
In contrast with Cubical Agda, Coq does not support quotient types at all. So how can we support proof repair across quotient type equivalences when quotient types do not even exist? The answer is to work with setoids—an external notion of quotients in Coq (Section 5.1). We extend the Pumpkin Pi proof term transformation to work with setoids (Section 5.2). Correctness yet again becomes an external notion we can only fully express metatheoretically, as it was in prior work in Coq (Section 5.3). A notable positive is that we are able to reuse and extend Pumpkin Pi’s existing proof automation to build our prototype automation (Section 5.4).
5.1. Quotient Types: External via Setoids
In Coq, we capture quotients using setoids, which are types paired with an equivalence relation representing equality (Sozeau, 2023). For example, we can represent our $\mathbb{N}/2$ type from Section 2.1 in Coq as the setoid (nat, mod_two), where mod_two is again equivalence modulo $2$. In this case, we call nat the carrier of the setoid. Unlike in Cubical Agda, there is no equivalence class constructor, and we instead define our functions and theorems on elements of the carrier.
Coq has a Setoid type class, and any setoid defines an instance of this type class. However, our transformation will not make use of instances of Setoid. Instead, we will understand metatheoretically that any pair of a type and an equivalence relation on that type forms a setoid, and use the machinery that derives from instances of the Equivalence and Proper type classes. Notably, any type forms a setoid with the equivalence relation being equality, so all types can be consider as setoids in this way.
When comparing elements of a setoid, the equivalence relation is used in place of equality. As a consequence of this, users of setoids need to juggle multiple notions of equality, unlike with native quotient types where the same equality is used universally. To use the running example, in Cubical Agda we compare equality of elements of $\mathbb{N}/2$ using the native equality $a\equiv b$, but in Coq, when considering (nat, mod_two) as a setoid, it is generally too strong to claim that a = b. Instead, we compare equality of elements of the setoid (nat, mod_two) by using the equivalence relation mod_two a b.
Coq’s setoids do not enforce that functions defined on them respect the equivalence relation until a user needs to do rewriting under that function in a proof, unlike with native quotient types, where well-definedness is checked statically upon writing the function. For example, we can write the function:
  f (x : nat) : bool := eqb x 3
for our (nat, mod_two) setoid, and Coq allows defining this function, since it is a valid function over the carrier of the setoid. However, unlike with equality, it is not the case that f x1 = f x2 whenever mod_two x1 x2, and we could not rewrite the term f x1 to f x2 using a proof that mod_two x1 x_2. When it is the case that, for setoids (A, eqA) and (B, eqB) and a function f : A -> B, the theorem:
  forall (a1 a2 : A), (eqA a1 a2) -> eqB (f a1) (f a2)
holds, we say that f is proper. A function satisfying this property defines an instance of the Proper type class in Coq. One example of a proper function is isEven : nat -> bool, which sends even numbers to true and odd numbers to false. This is proper considering the domain as the setoid (nat, mod_two) and the codomain as the setoid (bool, =).
To do repair on Cubical Agda, we had types that were strictly equivalent after taking quotients. In Coq, we instead have a notion of a setoid equivalence. Two setoids (A, eqA) and (B, eqB) are equivalent if there is a pair of functions f : A -> B and g : B -> A satisfying the following properties:
  • f and g are proper.
  • forall (a : A), eqA (g (f a)) a
  • forall (b : B), eqB (f (g b)) b
As an example, the function isEven defined above is one half of an equivalence between (nat, mod_two) and (bool, =). It has as an inverse the function sending true to 0 and false to 1. It is easy to verify that this pair of functions satisfies the above properties.
5.2. Transformation: Extended
To adapt our transformation to setoids, we adapt our transformation to handle the changes in how equality works (Figure 5). Our transformation in Coq currently supports only the case where the source type of the transformation uses the native equality and the target type is any setoid. Because equality is an inductive type, there are three ways it manifests in terms if no axioms are used: its type, its constructors, and its eliminators.
Equality Types
For any (a1 a2 : A), there is the native type @eq A a1 a2 representing proofs of equality between a1 and a2. Every equality in the source type is translated into an equivalence relation in the target type. We require that, for each type C which lifts to a type D, the user specify an equivalence relation equiv_D and a proof that relation is an equivalence relation. In Coq, such proofs are given as instances of the Equivalence type class. If the user does not provide an equivalence relation, native equality is used instead. Then, all occurrences of @eq C are replaced with equiv_D by the transformation, as seen in the EqApp rule. Notice that we do not transform instances of @eq which are not applied to a type; this is because, until @eq is applied to some type, the transformation cannot determine which equivalence relation it should be transformed into.
Constructors
Equality has one constructor,
  @eq_refl : (A : Type) -> (a : A) -> @eq A a a
We know that @eq C will be transformed into equiv_D, and thus we must produce a term of type forall (d : D), equiv_D d d. Because the user proves that their relations are equivalence relations, each carries a proof of that term, which we will denote by reflexivity D. We can therefore replace @eq_refl C with reflexivity D, as seen in the EqReflApp rule. Again, notice that we do not repair @eq_refl unless it is applied to a type. This is for the same reason as before; until @eq_refl is applied to a type, the transformation does not know which proof of reflexivity to transform it into.
Eliminators
In Coq, equality has three eliminators, each for different sorts, but we will only focus on the Type-sorted eliminator here, @eq_rect. Equality in Coq is Leibniz equality, meaning that for any P : C -> Type, if c1 = c2 then P c1 -> P c2. Thus, the eliminators define term rewrites, with P defining where in the term rewrites take place. Specifically, for any application:
  @eq_rect (A : Type) (x : A) (P : A -> Type) (px : P x) (y : A) (H : x = y)
[…]
flushleft $\Gamma$Uparrow $\Gamma^{\prime}$
[LiftEmpty]
[LiftCons] Γ⊢x ⇑x’
Γ⊢X ⇑X’
Γ⇑Γ’ (Γ, x : X) ⇑(Γ’, x’ : X’)
[…]
[EqReflApp] Γ⊢A ⇑B Γ⊢@eq_refl (A) ⇑reflexivity(B)
[EqRewrite] Γ⊢A ⇑B
Γ⇑Γ’
Γ⊢x ⇑x’
Γ⊢P ⇑P’
Γ⊢f ⇑f’
Γ⊢y ⇑y’
Γ⊢e ⇑e’
[…]
Figure 5. The additional rules needed for repairing to a setoid. There are two mutually defined judgements. The first defines lifting of environments, and the second defines lifting of terms.
5.3. Correctness: External
[…]
Note that, even though correctness criteria like elimOK cannot in general be stated and proven internally to Coq without axioms, specific instantiations of those criteria to specific functions and proofs are sometimes provable internally. This is especially true for functions without any dependent types, for which we solely need to show that equivalent inputs map to equivalent outputs. In general, however, internalizing univalence in even an ad hoc way in Coq requires at least functional extensionality (Tabareau et al., 2021). Alternatively, one could use the homotopy type theory library for Coq, but this would rely wholly on the univalence axiom (Bauer et al., 2017).
In any case, since the proof term transformation by definition results in a proof term, that proof term can always be checked by the type checker. This means the fact that correctness is not guaranteed is not in itself too worrying. If one is solely concerned with theorems holding, and not how they are proven, then the lack of internal correctness just means that it is up to the user to check that the repaired theorem statements are the same as the original theorem statements up to the change in datatype. When one cares about the contents of functions and proofs, it is up to the user to check that those contents are the same as the original up to the change in the datatype. But the proofs produced will either hold or, in the event of a mistake in the algorithm or its implementation, simply not make it past the type checker.
5.4. Automation: Prototype
The transformation we described can be implemented in Coq using the plugin system. We have implemented a prototype of this as an extension to the Pumpkin Pi plugin. Plugins in Coq are a method of adding additional functionality to Coq. They are written in OCaml, which is the metalanguage Coq is implemented in, and can interact directly with the Coq codebase. Plugins can directly trasform and produce terms; all terms that plugins produce are checked by Coq’s type checker, and so cannot be ill typed.
Pumpkin Pi is a plugin which automatically performs proof repair across type equivalences. Pumpkin Pi has various classes of proof repair transformations across type equivaleces for which it has specialized automation. We add an additional class, termed setoid lifting, to support our extended transformation. This class mostly reuses the existing transformation, but when a source term matching the conditions set in the EqApp, EqReflApp, and EqRewrite rules is encountered, the term is translated according to those rules. In all, our extension was 763 lines of code. The implementation of these lifting rules can be found in lift.ml, liftconfig.ml, and liftrules.ml in the supplementary material.
Unlike for EqApp and EqRefl, performing the EqRewrite rule is not done as a one-for-one term substitution. Recall that we assumed the existence of an oracle which could produce proofs of rewrites. To implement this oracle, we rely on Coq’s setoid automation. Coq has a tactic, called setoid_rewrite, which attempts to perform rewriting by an equivalence relation. However, because our equivalence relations are not generally Leibniz, we must prove for each function we define that the function is proper, as defined in Section 5.1, if we wish to rewrite under applications of that function. When we prove this statement, we instantiate the Proper type class. Furthermore, the proofs that our relations are equivalences are themselves instances of the Equivalence type class. The setoid_rewrite tactic uses these type class instances to search for proofs of rewrites, and thus we can use it as our oracle.
Unlike with applications of @eq_rect, however, we cannot directly specify specific locations where we wish to perform rewrites by providing a motive P : B -> Type. To get around this, if we are trying to prove the rewrite P x -> P y, we perform a substitution P[z/y], where z is free in P, and then define P := fun z => P[z/y]. Then, for another fresh variable w, we can use the setoid_rewrite tactic to prove forall (w : B), P w x -> P w y, and recover our desired rewrite proof as (forall (w : B), P w x -> P w y) y.
Theorem depRec (C : Type)
  (posP : forall (n : nat), C)
  (negSucP : forall (n : nat), C)
  (z : GZ) :
  C.
Theorem depElimProp (P : GZ -> Prop)
  ‘(p : Proper (GZ -> Prop) (eq_GZ ==> iff) P)
  (posP : forall (n : nat), P (depConstrPos n))
  (negSucP : forall (n : nat), P (depConstrNegSuc n))
  (z : GZ) :
  P z.
Figure 6. The types of the two eliminators we use in one of our case studies. The left has non-dependently typed output, but can eliminate into Type, while the right has dependently typed output but only eliminates into Prop. The right eliminator also requires a proof that the motive is proper as a function from the setoid (Z, eq_GZ) to the setoid (Prop, iff).
To facilitate using the setoid automation, we restrict the class of objects we can repair. Specifically, whereas the original proof repair work had a single dependent eliminator, we have multiple eliminators. One eliminates into the sort Type, but is purely nondependent; in the language of Cubical Agda, this would be called a recursor. The other is dependent, but only eliminates into the sort Prop, and requires that the output motive of the eliminator be proven to be proper, considering Prop as a setoid with the if-and-only-if equivalence relation. The types of two of these eliminators for one of our case studies are in Figure 6.
The reason for this split is to facilitate automation. The setoid automation works for equivalence relations on a fixed type. As such, we cannot instantiate an instance of Equivalence for a dependently typed notion of an equivalence relation. If our eliminator into Type was dependently typed, we would then not be able to use the setoid automation to perform rewrites on applications of functions we define. For our Prop-sorted eliminator, this loss means that users cannot automatically perform rewrites on the proofs of propositions. In Coq, Prop is frequently treated as effectively proof irrelevant (despite not actually being proof irrelevant without axioms), so this loss is more acceptable.
Note that, to perform repair, we presently require that users annotate their proofs prior to running repair. This is consistent with previous work using Pumpkin Pi, where annotations were required to identify parts of the configuration to the tool, thereby decoupling the undecidable part of proof repair (configuration inference) from the decidable part (the proof term transformation itself). Even in cases where general inference is undecidable, however, Pumpkin Pi can sometimes attempt to discover these components anyway using custom heuristics for unification for particular classes of changes. We have yet to implement such inference for this class of changes. Thus, users need to explicitly use the defined dependent constructors, dependent eliminators, and iota-reduction theorems when defining terms.
Like in previous work using Pumpkin Pi, we do not directly repair terms involving pattern matching and recursion, including on equality proofs. Pumpkin Pi includes some automation to transform pattern matching and recursion to induction, which comes bundled in our extended prototype. Finally, we presently require that all instances of @eq and @eq_refl are applied to a type, and that the eliminators for @eq are fully applied.
@@ 6. Case Studies
We apply our extended proof term transformations for two proof repair case studies that use quotient type equivalences. First, we conduct repair between two representations of the integers (Section 6.1). Second, we study two common implementations of the queue data structure, and how we can repair from one to the other (Section 6.2). We do each of these case studies in both Cubical Agda (manually) and Coq (with our prototype automation when possible), demonstrating the tradeoffs of the internal and external views directly on these case studies. All of the case study examples can be found in more detail in both Cubical Agda and Coq in supplementary material.
@@ 6.1. Adding, Fast and Slow
Our first case study is mathematically motivated. We consider a change in the type representing integers. We repair addition and proofs about addition from one representation to the other. Finally, we recover the repaired proofs for a more efficient version of addition over the repaired type. Our Cubical Agda and Coq proofs for this case study can be found in grothendieck_int_equivalence.agda and grothendieck_int_equivalence.v respectively.
@@
Inductive Z : Set :=
| pos : nat -> Z
| negsuc : nat -> Z.
Definition GZ := nat * nat.
Definition eq_GZ (z1 z2 : GZ) :=
  match z1, z2 with
  | (a1, a2), (b1, b2) => a1 + b2 = a2 + b1
  end.
Instance eq_GZ_equiv : Equivalence eq_GZ.
Instance GZ_setoid : Setoid GZ :=
  {equiv := eq_GZ ;
  setoid_equiv := eq_GZ_equiv}.
Figure 7. The types of our integer representations in both Cubical Agda (left) and Coq (right). The Setoid instance is not explicitly needed, but is included for reader clarity.
@@
We start with the default implementation of the integers in Cubical Agda—two copies of $\mathbb{N}$ glued together with one of them reversed, which we call $\mathbb{Z}$. We will repair functions and proofs about $\mathbb{Z}$ to use a representation of integers that may be more familiar to set theorists: Instead of two copies of $\mathbb{N}$ glued together at the ends, we consider the integers as elements of $\mathbb{N}\times\mathbb{N}/\sim$, where $(x_{1},x_{2})\sim(y_{1},y_{2})\iff x_{1}+y_{2}=x_{2}+y_{1}$. [^6] We can view this an an instance of constructing the Grothendieck group from the commutative monoid $\mathbb{N}$, and hence the integers arise as the unique group satisfying the universal property that any monoid homomorphism out of $\mathbb{N}$ can be uniquely extended to a group homomorphism out of $\mathbb{Z}$. We refer to the resulting quotient type as GZ.
The definitions of $\mathbb{Z}$ and GZ in both Cubical Agda (using quotient types) and Coq (using setoids) can be found in Figure 7. In Cubical Agda, we define the equivalence relation and then quotient $\mathbb{N}\times\mathbb{N}$ by that relation. As quotient types are internal to Cubical Agda, we are obligated to show that the relation is an equivalence relation in the construction of the quotient itself. In Coq, we still define the equivalence relation, but never formally take a quotient. Instead, we prove that the equivalence relation is an instance of the Equivalence type class in order to use Coq’s setoid rewriting automation.
In order to do repair between these two types, we need them to be isomorphic. In this case, the isomorphism is the map we expect: pos n maps to [(n , 0)], while negsuc n maps to [(0 , n + 1)]. Verifying that this map is bijective is straightforward in Cubical Agda. In Coq, because we never form equivalence classes and instead work over the carrier of the setoid, the map sends pos n to (n, 0) and negsuc n to (0 , n + 1). This map can also be seen to satisfy the definition of setoid equivalence given in Section 5.1.
@@
To repair functions and proofs across this change, we first must decompose our isomorphism into a configuration. The full configuration for Cubical Agda can be found in Figures 15 and  16 in the appendix, while the full configuration in Coq can be found in Figures 17 and  18 in the appendix. There, we can see the dependent constructors, dependent eliminators, and $\iota$-reduction rules for both types.
The types of the components of the configuration correspond with each other, and both sides of the configuration share the same inductive structure. In Cubical Agda, the only change in types here compared to Pumpkin Pi is that every motive P : GZ normal-$\to$ Set comes with the requirement that ((x : GZ) normal-$\to$ isSet (P x)). In Coq, the configuration differs from those found in Pumpkin Pi in that there are multiple eliminators. One eliminator, which we term depRec, can only eliminate into nondependent types. The other eliminator, which we term depElimProp, can eliminate into dependent types, but only those that reside in Prop. Furthermore, to use depElimProp on GZ, we must prove that the motive P : GZ -> Prop we supply to the eliminator is a proper function, where the sort Prop is viewed as the setoid (Prop, iff). We show these eliminators in Figure 6. In addition, each of these eliminators would their own need their own set of $\iota$-reduction theorems. However, needing to $\iota$-reduce an application of depElimProp in Coq will be rare in practice, because Prop is frequently informally treated as proof irrelevant. As such, we only provide the $\iota$ rules for depRec.
Thus, with the configuration defined, we can perform repair according to the procedures discussed in Sections 4.2 and 5.2. To give one concrete example of a term being repaired, we can consider the case of repairing addition from $\mathbb{Z}$ to GZ. First, we give the definition of addition on $\mathbb{Z}$ from the Cubical Agda standard library, as well as the same algorithm implemented in Coq in Figure 8. Notice that in Cubical Agda, we can use our general eliminator, while in Coq we use our recursor.
@@
Definition addZ (z1 z2 : Z) : Z :=
  depRecZ
    Z
    (fun (p : nat) => add_posZ z1 p)
    (fun (p : nat) => add_negsucZ z1 p)
    z2.
Figure 8. The standard definition of addition on $\mathbb{Z}$ in Cubical Agda (left) and the corresponding Coq implementation (right), explicitly annotated with the dependent eliminators.